The csfax Class for L^{A}TEX 2ε

Martin Reddy

February 7, 1997

1 Introduction

This document describes the csfax class (v1.3) which has been written for L^{A}TEX 2ε. This class provides a generic facsimile transmission sheet for use in the department of Computer Science. It exports the macro \makefax which generates a fax cover sheet, along with a number of other macros for specifying the details on the cover sheet. You can use this class by simply starting your L^{A}TEX file with:

     \documentclass{csfax}

2 Cover Page Information

The \makefax command produces a fax cover page. You should place this at the start of your document so that the cover sheet appears on the first page. You can use the following macros to define the various details which are displayed on the fax transmission sheet:

• \FaxFromName{<name>} : the name of the sender of the fax (i.e. probably you). You should always specify this field.

• \FaxFromEmail{<email>} : the full e-mail address for the sender of the fax (i.e. probably your e-mail).

• \FaxFromExt{<ext>} : the (four figure) extension number for your office. E.g. if your internal extension number is 501234, then you should just specify the last four digits, i.e. 1234.

• \FaxShowFromFax : if you include this command in your document then csfax will generate an entry with the departmental fax number along after your other personal information such as name, telephone number, etc. Note that this is already given in the address header, but you may wish to replicate the information in the From: section for thoroughness.

• \FaxIntCode : by default, all telephone numbers will be provided in UK format (i.e. 0131 STD code). If you specify this command, then all numbers will be given in international format (i.e. +44 131).

• \FaxToName{<name>} : the name of the intended recipient of the fax. You should always specify this field.

• \FaxToOrg{<org>} : the name of the organisation or department where the recipient works. This may be useful to help the destination fax operator locate the recipient.

• \FaxToTel{<telno>} : the telephone number of the recipient, if known.

• \FaxToFax{<faxno>} : the fax number of the recipient, if known.

• \FaxMessage{<text>} : any message that you wish to appear on the fax cover sheet.

• \FaxDate{<date>} : the date of transmission. If you do not specify this, then the current date will be assumed.

• \FaxPages{<number>} : the number of pages being transmitted (including the cover sheet). If you do not specify this, then sufficient space will be left on the fax cover page for you to write in the page count manually.

3 The MakeFax Script

Often you will only want to produce a one-off fax cover sheet and so you may not wish to go to all the lengths of writing and compiling a L^{A}TEX document just for this. A utility has therefore been provided to do all this for you: makefax.

The makefax utility will prompt you for the various pieces of information and will then create a L^{A}TEX document and compile it behind the scenes. You are then given the option to output this directly to a printer, or to save the DVI file in the current directory. All temporary files are removed upon exit.

The program may also be customised via a number of command line options which allow you to explicitly specify any field. Default values for personal fields will be taken from your ph database entry.

Type makefax -h for a brief summary of the command line options; or refer to man makefax for a more detailed discussion.

A Example L^{A}TEX 2ε Code

%
% An example of how to use the csfax class to
% produce a facsimile cover sheet.
% Version 1.3. M. Reddy (mxr), 7 Feb 1997.
%
\documentclass{csfax}
\FaxIntCode
\FaxShowFromFax
\FaxFromName{Martin Reddy}
\FaxFromEmail{M.Reddy@ed.ac.uk}
\FaxFromExt{5164}
\FaxToName{Bill Gates}
\FaxToOrg{Microsoft Corp.}
\FaxToTel{don't know}
\FaxToFax{don't know either}
\FaxPages{1}
\FaxMessage{%
  Hi there Bill!
  Could you lend us a couple million mate?
}
\begin{document}
\makefax
\end{document}

      Above L^{A}TEX source available on-line in /usr/local/share/TeX/tex/latex/examples